OpenStudy (anonymous):

A car is traveling at a rate of 36 miles per hour. What is the car's rate in feet per second? How many feet will the car travel in 2 seconds? In your computations, use the fact that 1 mile is equal to 5280 feet. Do not round your answers.

OpenStudy (anonymous):

@ehuman @phi @mathmale

OpenStudy (phi):

can you change 36 miles to feet ?

OpenStudy (anonymous):

i'm not sure

OpenStudy (phi):

you want to multiply \[ \text{miles} \cdot \frac{\text{feet}}{\text{miles}} \] in other words you need the ratio of feet per mile. Can you write down the ratio of feet per mile?

OpenStudy (mathmale):

We need to convert 36 miles per hour to feet per sec (ft/sec). We know two crucial facts: 1hr=3600 sec., and 1 mile = 5280 feet. 1) convert miles to feet: \[\frac{ 36 miles }{ hour }*\frac{ 5280 feet }{ 1 mile }\]

OpenStudy (mathmale):

(2) Convert hours to seconds: \[\frac{ 36 miles }{ 1 hour }\frac{ 5280 feet }{ 1 mile }\frac{ 1 hour }{ 3600 \sec }.\] Cancel "hour" and "miles". Multiply as indicated.

OpenStudy (phi):

1 mile is equal to 5280 feet you use that to write 1 mile per 5280 feet or \[ \frac{ \text{1 mile}}{\text{5280 feet}}\] but that is upside down (we want feet "up top") so "flip it" \[ \frac{\text{5280 feet}}{ \text{1 mile}}\] that is the ratio you use to convert miles to feet \[ 36 \text{ miles} \cdot \frac{\text{5280 feet}}{ \text{1 mile}} \] will give you the number of feet in 36 miles.